A charge of 20 x 10-6 C is 20 cm from another charge of 60 x 10-6 C. (a) Find the force between them. (b) Determine the Ep midway between them.

1
Expert's answer
2022-02-24T09:04:47-0500

a)

"F=\frac{k(q_1^2-q_2^2)}{r^2}=720~N,"

b)

"E=\frac{4k(q_1-q_2)}{r^2}=36~\frac{MV}m."